OpenStudy (anonymous):

A company is to paint 3000 metre in 30 days. The boss hires 50 men and they work 8hrs per day. After 20 days, 1200m is completed. How many more men does the boss have to employ to complete on time if each man work 10 hours per day ?

OpenStudy (anonymous):

@One098

OpenStudy (one098):

@sleepyjess

OpenStudy (one098):

@paki

OpenStudy (sleepyjess):

Mk, I'm gonna try this, but if it doesn't work out work out, I will tag someone else :)

OpenStudy (sleepyjess):

First, I'm thinking to find the number of meters a day, then number an hour?

OpenStudy (paki):

tell me... its a direct proportion or inverse....?

OpenStudy (anonymous):

excuse me @HatcrewS do you by any chance can read and speak chinese

OpenStudy (anonymous):

@paki i dont know, is it direct ?

OpenStudy (anonymous):

Hello.

OpenStudy (anonymous):

hi there

OpenStudy (sleepyjess):

@OpiGeode , if you are not going to contribute to the question, please do not spam this post.

OpenStudy (mathmath333):

if more men work the number of days the work would complete early if the more hour they work daily the work would complete early so \(m~~\alpha\dfrac{1}{w}\) \(h ~~\alpha\dfrac{1}{w}\) so \(h \times m~~\alpha\dfrac{1}{w}\) hence \(h \times m~~=\dfrac{k}{w}\) \(k~~=8\times 50\times 1200\) now they have done work of \(1200\) the work need to be done is \(3000-1200=1800\) so \(h \times m~~=\dfrac{k}{w}\) \(10 \times m~~=\dfrac{8\times 50\times 1200}{1800}\) find \(m\)

OpenStudy (anonymous):

thank you sir @mathmath333